OpenStudy (anonymous):

Which of the following ratios correctly describes the cosine function? A. opp/adj B. app/hyp C. adj/hyp

OpenStudy (anonymous):

b

OpenStudy (anonymous):

SOH-CAH-TOA

OpenStudy (anonymous):

The answer was C.

OpenStudy (anonymous):

Do not forget in trigonomatry SOH= Sin=Opp/Adj CAH=Cosine=Adj/Hyp TOA=Opp/Adj

OpenStudy (anonymous):

SORRY TYPING MISTAKE